40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 31 17:11:40 2012, jamie, Update, Alignment, progress, then setback 
    Reply  Wed Oct 31 17:18:34 2012, Jenne, Update, Alignment, progress, then setback - trend of BS table shift PRM_BS_table_bumped_ITMY_MC1_no_big_change_2hoursBack.png
       Reply  Wed Oct 31 17:36:39 2012, jamie, Update, Alignment, progress, then setback - trend of BS table shift 
       Reply  Thu Nov 1 01:51:37 2012, rana, Update, Alignment, progress, then setback - trend of BS table shift 
          Reply  Thu Nov 1 08:48:42 2012, steve, Update, Alignment, progress, then setback - trend of BS table shift alignmentstrugle2d.pngalignmentstrugle2da.pngbefpd09252012.png
       Reply  Thu Nov 1 10:58:49 2012, jamie, Update, Alignment, progress, then setback - trend of BS table shift foo.png
Message ID: 7646     Entry time: Wed Oct 31 17:11:40 2012     Reply to this: 7647
Author: jamie 
Type: Update 
Category: Alignment 
Subject: progress, then setback 

jamie, nic, jenne, den, raji, manasa

We were doing pretty well with alignment, until I apparently fucked things up.

We were approaching the arm alignment on two fronts, looking for retro-reflection from both the ITMs and the ETMs.

Nic and Raji were looking for the reflected beam off of ETMY, at the ETMY chamber.  We put an AWG sine excitation into ETMY pitch and yaw.  Nic eventually found the reflected beam, and they adjusted ETMY for retro-reflection.

Meanwhile, Jenne and I adjusted ITMY to get the MICH Y arm beam retro-reflecting to BS.

Jenne and I then moved to the X arm.  We adjusted BS to center on ITMX, then we moved to ETMX to center the beam there.  We didn't both looking for the ETMX reflected beam.  We then went back to BS and adjusted ITMX to get the MICH X arm beam retro-reflected to the BS.

At this point we were fairly confident that we had the PRC, MICH, and X and Y arm alignment ok.

We then moved on the signal recycling cavity.  Having removed and reinstalled the SRC tip-tilts, and realigning everything else, they were not in the correct spot.  The beam was off-center in yaw on SR3, and the SR3 reflected beam was hitting low and to the right on SR2.  I went to loosen SR3 so that I could adjust it's position and yaw, and that when things went wrong.

Apparently I hit something BS table and completely lost the input pointing.  I was completely perplexed until I found that the PZT2 mount looked strange.  The upper adjustment screw appeared to have no range.  Looking closer I realized that we somehow lost the gimble ball between the screw and the mount.  Apparently I somehow hit PZT2 hard enough to separate from the mirror mount from the frame which caused the gimble ball to drop out.  The gimble ball probably got lost in a table hole, so we found a similar mount from which we stole a replacement ball.

However, after putting PZT2 back together things didn't come back to the right place.  We were somehow high going through PRM, so we couldn't retro-reflect from ITMY without completely clipping on the PRM/BS apertures.  wtf.

Jenne looked at some trends and we saw a big jump in the BS/PRM osems.  Clearly I must have hit the table/PZT2 pretty hard, enough to actually kick the table.  I'm completely perplexed how I could have hit it so hard and not really realized it.

Anyway, we stopped at this point, to keep me from punching a hole in the wall.  We will re-asses the situation in the morning.  Hopefully the BS table will have relaxed back to it's original position by then.

ELOG V3.1.3-